You are concerned that nausea may be a side effect of Tamiflu, but you cannot just give Tamiflu to patients with the flu and say that nausea is a side effect if people become nauseous. This is because nausea is common for people who have the flu. From past studies you know that about 30% of people who get the flu experience nausea. You collected data on 1595 patients who were taking Tamiflu to relieve symtoms of the flu, and found that 525 experienced nausea. Use a 0.01 significance level to test the claim that the percentage of people who take Tamiflu for the relief of flu symtoms and experience nausea is greater than 30%.
Calculate your test statistic. Write the result below, and be sure to round your final answer to two decimal places.
Calculate your p-value. Write the result below, and be sure to round your final answer to four decimal places.

Answer by Boreal(15235)

Point estimate is p hat=525/1595=0.32915
Ho: p<=0.30
Ha: p>0.30
alpha=0.01 p{reject Ho|Ho true}
test statistic is a z
critical value is z>2.326
z=(p hat-p)/sqrt(p*(1-p)/n)
=0.02915/sqrt( (0.3*0.7/1595); sqrt term =0.01147
z=2.54
reject Ho
p-value z>2.540 which is 0.0055.